# Tropicalization of curve arrangement complements and arroids

### v2

@@ 1. Introduction
When all the multiplicity functions of an arroid are constant taking value one, the arroid is said to be transversal. We construct a fan associated to each transversal arroid, and the following theorem shows that such a fan is a tropical variety, i.e. satisfies the balancing condition of tropical geometry, see e.g. [8, 22] for definitions. {introthmTemp:1} For each transversal arroid $\mathbf{A}$, there is a fan $\Sigma_{\mathbf{A}}$, called the fan of $\mathbf{A}$, which is a balanced tropical variety. Using transversal arroids, we proceed to study the tropicalization of the complements of certain types of arrangements of curves. An arrangement of curves is very affine if it contains at least three lines intersecting generically, and transverse if all curves of the arrangement intersect pairwise transversely. In Section 4.4, we show that for a transverse very affine arrangement of curves, the tropicalization of the complement is computed by the arroid fan. {introthmTemp:2} Let $\mathcal{B}$ be a transverse very affine arrangement of curves in the plane $\mathbb{P}_{K}^{2}$. Then the tropicalization $\operatorname{trop}(X_{\mathcal{B}})$ of the complement is supported on the fan of the associated transversal arroid $\mathbf{A}_{\mathcal{B}}$. For arrangements of lines, Theorem 4.15 recovers that the tropicalization of the complement is computed using the rank three matroid of the arrangement (see e.g. [22, Theorem 4.1.11]), using the Ardila–Klivans fan structure [7]. The difficulty in generalizing beyond the transverse case lies primarily in understanding the resolution of singularities that arise when higher order intersections are allowed in the arrangement, reflecting an earlier observation found in [10, p.20].
@@
Next, we turn to relating the cohomology of the complement of a very affine transverse arrangement of curves to the tropical cohomology of the fan of its associated transversal arroid. For a reminder on tropical cohomology, see Section 2.4. For line arrangements and their corresponding matroids, an isomorphism between the matroid Orlik–Solomon algebra [25], computing cohomology of the complement using only the intersection properties recorded by the matroid, and tropical cohomology of the matroid fan, was described by Zharkov [32]. We consider transverse very affine arrangements $\mathcal{B}$ of non-singular rational curves in $\mathbb{P}_{\mathbb{C}}^{2}$, i.e. of lines and conics, such that that no intersection point of the arrangement contains exactly the same curves. Such an arrangement will be called simple.
@@
In Proposition 6.1, we show that the complement of a simple arrangement is wunderschön in the sense of [2, Definition 1.2], which is in this context primarily a restriction on its mixed Hodge structure. In light of [2, Theorem 6.1], this implies that the complement of simple arrangements are cohomologically tropical i.e. its rational cohomology can be computed using the $\mathbb{Q}$-coefficient tropical cohomology of its tropicalization, if and only if the corresponding arroid fan is a tropical homology manifold. {introthmTemp:3} Let $X_{\mathcal{B}}$ be the complement of a simple arrangement $\mathcal{B}$. Then $X_{\mathcal{B}}$ is cohomologically tropical if and only if the corresponding arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$ is uniquely balanced along each of its rays. This result follows from equivalent conditions for an arroid fan to be a tropical homology manifold given in Theorem 5.4, and we study which conditions this imposes on curve arrangements in Section 5.2.
@@
Using Theorem 6.2, we study the question of maximality for a real arrangement and its complexification. Let $X$ be a complex variety defined over $\mathbb{R}$, with $X(\mathbb{R})$ its set of real points and $X(\mathbb{C})$ its set of complex points. The Smith-Thom inequality gives bounds for the sum of the $\mathbb{Z}/2\mathbb{Z}$-Betti numbers as follows,
@@
In light of these conditions, in Theorem 7.3 we give conditions for the complement of a simple arrangement of real curves in the plane to be maximal using its tropicalization. Moreover, using the wunderschön and cohomologically tropical properties, we give the following concrete construction of varieties satisfying the above conditions. {introthmTemp:4} Let $\mathcal{B}$ be a simple arrangement of real curves in $\mathbb{P}_{\mathbb{C}}^{2}$, with all intersection points being real, and such that the tropicalization $\operatorname{Trop}(X_{\mathcal{B}})$, which is supported on the arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$, is a tropical homology manifold. Then the following four properties are satisfied:
@@
This theorem is illustrated by providing an infinite family of maximal surfaces in Example 7.4, which give examples of the types of variety required in [4], using conditions (a), (b) and (c). Condition (d) can be compared to the bounds for Betti numbers given in [27].
The paper is structured as follows. In Section 2, we recall the notions of geometric tropicalization, tropical modifications, and tropical cohomology. In Section 3, we illustrate how geometric tropicalization can be used to compute the tropicalization of the complement of an arrangement of curves in the plane. In Section 4, we introduce arroids as an abstract generalization of arrangements of curves, define fans associated to arroids which are transversal (see Definition 4.9), and relate the tropicalization of the complement of an arrangement to the fan of the associated arroid in the transversal case. Next, in Section 5, we study which arroid fans are tropical homology manifolds, and give some conditions for the arroid fan of an arrangement of lines and conics to be a tropical homology manifold. Finally, in Section 7, use the concepts developed in the rest of the paper to study the maximality of arrangements of lines and conics.
@@ 2.2. Geometric tropicalization
We briefly recall the geometric tropicalization approach to tropicalization. Let $\mathbf{T}$ be an algebraic torus over the field $K$, with $M$ its lattice of characters and $N\coloneqq\operatorname{Hom}_{\mathbb{Z}}(M,\mathbb{Z})$ its dual lattice. Let $X\subset\mathbf{T}$ be a subvariety and $X^{\prime}$ a normal, $\mathbb{Q}$-factorial variety birational to $X$. Any divisor $D$ on $X^{\prime}$ induces a valuation $\operatorname{val}_{D}\colon K(X)\to\mathbb{Z}$ on the field of fractions $K(X)$ of $X$. Such a valuation is called a divisorial valuation on $K(X)$, and it defines a vector $[\operatorname{val}_{D}]\in N_{\mathbb{Q}}$ by restricting a character $m\in M$ to a rational function $m|_{X}\in K(X)$ and evaluating it using $\operatorname{val}_{D}$.
@@
Let $\Delta_{\partial X}$ be the boundary complex of $({\overline{X}},\partial X)$, i.e. a simplicial complex with vertices $\{1,\dots,m\}$ containing the simplex $\{i_{1},\dots,i_{k}\}$ if and only if $D_{i_{1}}\cap\dots\cap D_{i_{k}}$ is non-empty. The $m$ divisorial valuations $[\operatorname{val}_{D_{1}}],\dots,[\operatorname{val}_{D_{m}}]$ give rays in $N_{\mathbb{R}}$, and for each simplex $\sigma$ of $\Delta$ we form the cone $[\sigma]\coloneqq\mathrm{cone}([\operatorname{val}_{D_{i}}]\mid i\in\sigma)$ in $N_{\mathbb{R}}$.
@@ 2.4. Tropical (co)homology
Let $\Sigma$ be a $d$-dimensional fan in $N$, with unique minimal 00-dimensional cone denoted $\mathbf{0}$. Recall from Section 2.3 that $L_{\mathbb{Z}}(\sigma)\subseteq N$ is the (saturated) lattice parallel to $\sigma$. For each $p=1,\dots,d$, the $p$-th multi-tangent space is the vector subspace
@@
where $\mathbf{0}$ is the unique minimal 00-dimensional cone of $\Sigma$, for $p=0,\dots,d$. Similarly, the tropical homology groups defined by $H_{p,0}(\Sigma)=\mathcal{F}_{p}(\mathbf{0})$.
Equip each cone $\sigma$ with an orientation, and for $\tau$ a face of $\sigma$, let $\mathrm{sign}(\tau,\sigma)$ be $1$ if the chosen orientations of $\tau$ and $\sigma$ are compatible, and $-1$ otherwise. For each $p=1,\dots,d$, we may define the $p$-th tropical Borel–Moore complex $C_{p,\bullet}^{BM}(\Sigma)$ as follows
@@
The differential $\partial_{k}$ is defined as the sum of its components $(\partial_{k})_{\gamma,\delta}\colon\mathcal{F}_{p}(\gamma)\to\mathcal{F}_{p}(\delta)$, which is given by $\mathrm{sign}(\delta,\gamma)\cdot\iota_{\delta\prec\gamma}$ if $\delta$ is a face of $\gamma$, and 00 otherwise.
@@
For any tropical fan $\Sigma$ of dimension $d$, the weight function $\omega$ gives rise to a fundamental class $[\Sigma,\omega]\in H_{d,d}^{BM}(\Sigma)$, and there are cap product maps $\frown[\Sigma,\omega]\colon H^{p,q}(\Sigma)\to H_{d-p,d-q}^{BM}(\Sigma)$.
@@
Let $\Sigma$ be a tropical fan with fundamental class $[\Sigma,\omega]\in H_{d,d}^{BM}(\Sigma)$.
@@ 3. Tropicalizing complements of arrangements of curves
Using the above procedure, by Theorem 2.3 the tropicalization $\operatorname{trop}(X)$ can be equipped with a fan structure consisting of one ray for each element $D$ of $\mathcal{B}$, and one two-dimensional cone between the rays for $D$ and $D^{\prime}$ if these two curves intersect in $S$. Moreover note that one may ignore the restriction on the characteristic of $K$ by Remark 2.4. Therefore, to complete the description of the tropicalization $\operatorname{trop}(X)$, it remains to find the directions of these rays.
@@
[…]
Figure 1. Curve arrangement from Example 3.1 and the compactifying divisor.
@@
Consider the arrangement of curves consisting of the coordinate axes $L_{1},L_{2},L_{3}\subset\mathbb{P}_{K}^{2}$, as well as the conic $C$ passing through the three points of intersection $L_{i}\cap L_{j}$, as shown in Figure 1. The complement of the arrangement $X\coloneqq\mathbb{P}_{K}^{2}\setminus(L_{1}\cup L_{2}\cup L_{3}\cup C)$ is not compactified by a simple normal crossing divisor in $\mathbb{P}_{K}^{2}$, so we must blow up the three points $L_{i}\cap L_{j}$. This yields a configuration including three exceptional divisors $E_{1}$, $E_{2}$ and $E_{3}$, also displayed in Figure 1.
@@
[…]
@@ 4. Axioms for abstract arrangements of curves
In light of the description of the tropicalization of an arrangement of curves in Section 3, we now propose an axiomatization generalizing arrangements of curves. We show that for a given arrangement, the axiomatization data extracted from the arrangement is sufficient to compute its tropicalization in its intrinsic torus.
@@ 4.1. Axioms
By the above construction, any matroid has a corresponding arroid. In particular, nonrealizable matroids are not realizable as arroids either, when requiring the specified curves to satisfy the conditions imposed by the degree function and multiplicity functions, i.e. that they must be lines intersecting in one point. However, when removing these restrictions, it may be possible to find curve arrangements having the intersection properties of the original unrealizable matroids. For nonrealizability in higher degrees, see Example 4.6.
@@
We return to Example 3.1. Let $\mathcal{A}=\{1,2,3,4\}$, where $i=1,2,3$ corresponds to the line $L_{i}$, and $4$ corresponds to the conic $C$. The degree function $d\colon\mathcal{A}\to\mathbb{Z}_{>0}$ is given by $d_{i}=1$ for $i<4$ and $d_{4}=2$. The multiset of points is $\mathcal{P}=\{\{1,2,4\},\{1,3,4\},\{2,3,4\}\},$ corresponding to the set of intersection points of the lines and conic. The Bézout property of $\mathcal{P}$ follows directly from the Bézout theorem for $\mathbb{P}_{K}^{2}$.
[…]
Figure 2. Curve arrangement from Example 4.4.
@@
Consider now the generic arrangement of three lines and a conic displayed in Figure 2. As before, let $\mathcal{A}=\{1,2,3,4\}$, where $i=1,2,3$ corresponds to the line $L_{i}$, and $4$ corresponds to the conic $C$, with the degree function $d\colon\mathcal{A}\to\mathbb{Z}_{>0}$ given by $d_{i}=1$ for $i<4$ and $d_{4}=2$. The multiset structure of the points is now more evident as
@@
In general, given an arrangement of curves =$\mathcal{B}=\{C_{1},\dots,C_{n}\}$ on $\mathbb{P}_{K}^{2}$, we can create an arroid of rank two as follows. Let $\mathcal{A}\coloneqq\{1,\dots,n\}$ and the function $d$ takes $i$ to the degree of the curve $d_{i}$. For any subset $\mathbf{p}\subseteq\mathcal{B}$ such that $C_{I}=\cap_{i\in I}\;C_{i}\neq\emptyset$, let $\mathbf{p}_{I_{1}},\dots,\mathbf{p}_{I_{k}}$ be the components of $C_{I}$. For each $\mathbf{p}_{I_{j}}$, we include $\mathbf{p}$ as an element in $\mathcal{P}$ with multiplicity function $m_{\mathbf{p}}$ given by the intersection multiplicities of the curves at the component $\mathbf{p}_{I_{j}}$. Thus the point $\mathbf{p}$ may appear multiple times with potentially different multiplicity functions. Summarizing, we have show the following proposition.
@@
Figure 3. Diagrammatic representation of the nonrealizable arroid from Example 4.6, with the omitted line in dashes.
@@
Let $\mathbf{A}$ be the arroid encoding parts of the data of a (self-crossing) hexagon inscribed in a conic. Its data, most easily intuited from Figure 3, is given by the ground set $\mathcal{A}\coloneqq\{C,1,\dots,6\}$, with $d(C)=2$ and $d(i)=1$ for all $i$, the multiplicity functions are all constant taking value $1$, with multiset of points $\mathcal{P}$ consists of the points named $\mathbf{a}=\{1,3\}$, $\mathbf{b}=\{2,5\}$, $\mathbf{c}=\{4,6\}$, as well as the points $\{1,2,C\}$, $\{3,4,C\}$, $\{5,6,C\}$, $\{3,5,C\}$, $\{1,6,C\}$, $\{2,4,C\}$, $\{2,3\}$, $\{4,5\}$,$\{1,5\}$, and $\{2,6\}$.
Pascal’s theorem, also known as the hexagrammum mysticum theorem (the reader may prefer hexagrammum nonmysticum as the name of the above arroid), states that for curves satisfying the above data, the three points $\mathbf{a}$, $\mathbf{b}$ and $\mathbf{c}$ must be colinear. By omitting this line from the arroid, it can no longer be realized in the projective plane. This example also generalizes the non-Pappus matroid, in light of the generalization of Pappus theorem by Pascal’s theorem, and one might construct further nonrealizable examples by working with the Cayley–Bacharach theorem.

### v3

@@ Abstract.
This is the accepted version of the article published in Mathematica Scandinavica 131 (2025), no. 2. DOI: 10.7146/math.scand.a-156674
@@ 1. Introduction
When all the multiplicity functions of an arroid are constant taking value one, the arroid is said to be transversal. We construct a fan associated to each transversal arroid, and the following theorem shows that such a fan is a tropical variety, i.e. satisfies the balancing condition of tropical geometry, see e.g. [8, 22] for definitions. {introthmTemp:1} For each transversal arroid $\mathbf{A}$, there is a fan $\Sigma_{\mathbf{A}}$, called the fan of $\mathbf{A}$, which is a balanced tropical variety. Using transversal arroids, we proceed to study the tropicalization of the complements of certain types of arrangements of curves. An arrangement of curves is very affine if it contains at least three lines intersecting generically, and transverse if all curves of the arrangement intersect pairwise transversely. In Section˜4.4, we show that for a transverse very affine arrangement of curves, the tropicalization of the complement is computed by the arroid fan. {introthmTemp:2} Let $\mathcal{B}$ be a transverse very affine arrangement of curves in the plane $\mathbb{P}_{K}^{2}$. Then the tropicalization $\operatorname{trop}(X_{\mathcal{B}})$ of the complement is supported on the fan of the associated transversal arroid $\mathbf{A}_{\mathcal{B}}$. For arrangements of lines, Theorem˜4.15 recovers that the tropicalization of the complement is computed using the rank three matroid of the arrangement (see e.g. [22, Theorem 4.1.11]), using the Ardila–Klivans fan structure [7]. The difficulty in generalizing beyond the transverse case lies primarily in understanding the resolution of singularities that arise when higher order intersections are allowed in the arrangement, reflecting an earlier observation found in [10, p.20].
@@
Next, we turn to relating the cohomology of the complement of a very affine transverse arrangement of curves to the tropical cohomology of the fan of its associated transversal arroid. For a reminder on tropical cohomology, see Section˜2.4. For line arrangements and their corresponding matroids, an isomorphism between the matroid Orlik–Solomon algebra [25], computing cohomology of the complement using only the intersection properties recorded by the matroid, and tropical cohomology of the matroid fan, was described by Zharkov [32]. We consider transverse very affine arrangements $\mathcal{B}$ of non-singular rational curves in $\mathbb{P}_{\mathbb{C}}^{2}$, i.e. of lines and conics, such that that no intersection point of the arrangement contains exactly the same curves. Such an arrangement will be called simple.
@@
In Proposition˜6.1, we show that the complement of a simple arrangement is wunderschön in the sense of [2, Definition 1.2], which is in this context primarily a restriction on its mixed Hodge structure. In light of [2, Theorem 6.1], this implies that the complement of simple arrangements are cohomologically tropical i.e. its rational cohomology can be computed using the $\mathbb{Q}$-coefficient tropical cohomology of its tropicalization, if and only if the corresponding arroid fan is a tropical homology manifold. {introthmTemp:3} Let $X_{\mathcal{B}}$ be the complement of a simple arrangement $\mathcal{B}$. Then $X_{\mathcal{B}}$ is cohomologically tropical if and only if the corresponding arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$ is uniquely balanced along each of its rays. This result follows from equivalent conditions for an arroid fan to be a tropical homology manifold given in Theorem˜5.4, and we study which conditions this imposes on curve arrangements in Section˜5.2.
@@
Using Theorem˜6.2, we study the question of maximality for a real arrangement and its complexification. Let $X$ be a complex variety defined over $\mathbb{R}$, with $X(\mathbb{R})$ its set of real points and $X(\mathbb{C})$ its set of complex points. The Smith-Thom inequality gives bounds for the sum of the $\mathbb{Z}/2\mathbb{Z}$-Betti numbers as follows,
@@
In light of these conditions, in Theorem˜7.3 we give conditions for the complement of a simple arrangement of real curves in the plane to be maximal using its tropicalization. Moreover, using the wunderschön and cohomologically tropical properties, we give the following concrete construction of varieties satisfying the above conditions. {introthmTemp:4} Let $\mathcal{B}$ be a simple arrangement of real curves in $\mathbb{P}_{\mathbb{C}}^{2}$, with all intersection points being real, and such that the tropicalization $\operatorname{Trop}(X_{\mathcal{B}})$, which is supported on the arroid fan $\Sigma_{\mathbf{A}_{\mathcal{B}}}$, is a tropical homology manifold. Then the following four properties are satisfied:
@@
This theorem is illustrated by providing an infinite family of maximal surfaces in Example˜7.4, which give examples of the types of variety required in [4], using conditions (a), (b) and (c). Condition (d) can be compared to the bounds for Betti numbers given in [27].
The paper is structured as follows. In Section˜2, we recall the notions of geometric tropicalization, tropical modifications, and tropical cohomology. In Section˜3, we illustrate how geometric tropicalization can be used to compute the tropicalization of the complement of an arrangement of curves in the plane. In Section˜4, we introduce arroids as an abstract generalization of arrangements of curves, define fans associated to arroids which are transversal (see Definition˜4.9), and relate the tropicalization of the complement of an arrangement to the fan of the associated arroid in the transversal case. Next, in Section˜5, we study which arroid fans are tropical homology manifolds, and give some conditions for the arroid fan of an arrangement of lines and conics to be a tropical homology manifold. Finally, in Section˜7, use the concepts developed in the rest of the paper to study the maximality of arrangements of lines and conics.
@@ 2.2. Geometric tropicalization
We briefly recall the geometric tropicalization approach to tropicalization. Let $\mathbf{T}$ be an algebraic torus over the field $K$, with $M$ its lattice of characters and $N\coloneqq\operatorname{Hom}_{\mathbb{Z}}(M,\mathbb{Z})$ its dual lattice. Let $X\subset\mathbf{T}$ be a subvariety and $X^{\prime}$ a normal, $\mathbb{Q}$-factorial variety birational to $X$. Any divisor $D$ on $X^{\prime}$ induces a valuation $\operatorname{val}_{D}\colon K(X)\to\mathbb{Z}$ on the field of fractions $K(X)$ of $X$. Such a valuation is called a divisorial valuation on $K(X)$, and it defines a vector $\operatorname{val}_{D}]\in N_{\mathbb{Q}}$ by restricting a character $m\in M$ to a rational function $m|_{X}\in K(X)$ and evaluating it using $\operatorname{val}_{D}$.
@@
Let $\Delta_{\partial X}$ be the boundary complex of $({\overline{X}},\partial X)$, i.e. a simplicial complex with vertices $\{1,\dots,m\}$ containing the simplex $\{i_{1},\dots,i_{k}\}$ if and only if $D_{i_{1}}\cap\dots\cap D_{i_{k}}$ is non-empty. The $m$ divisorial valuations $\operatorname{val}_{D_{1}}],\dots,[\operatorname{val}_{D_{m}}]$ give rays in $N_{\mathbb{R}}$, and for each simplex $\sigma$ of $\Delta$ we form the cone $\sigma]\coloneqq\mathrm{cone}([\operatorname{val}_{D_{i}}]\mid i\in\sigma)$ in $N_{\mathbb{R}}$.
@@ 2.4. Tropical (co)homology
Let $\Sigma$ be a $d$-dimensional fan in $N$, with unique minimal 0-dimensional cone denoted $\mathbf{0}$. Recall from Section˜2.3 that $L_{\mathbb{Z}}(\sigma)\subseteq N$ is the (saturated) lattice parallel to $\sigma$. For each $p=1,\dots,d$, the $p$-th multi-tangent space is the vector subspace
@@
where $\mathbf{0}$ is the unique minimal 0-dimensional cone of $\Sigma$, for $p=0,\dots,d$. Similarly, the tropical homology groups defined by $H_{p,0}(\Sigma)=\mathcal{F}_{p}(\mathbf{0})$.
Equip each cone $\sigma$ with an orientation, and for $\tau$ a face of $\sigma$, let $\mathrm{sign}(\tau,\sigma)$ be $1$ if the chosen orientations of $\tau$ and $\sigma$ are compatible, and 1-1 otherwise. For each $p=1,\dots,d$, we may define the $p$-th tropical Borel–Moore complex $C_{p,\bullet}^{BM}(\Sigma)$ as follows
@@
The differential $\partial_{k}$ is defined as the sum of its components $(\partial_{k})_{\gamma,\delta}\colon\mathcal{F}_{p}(\gamma)\to\mathcal{F}_{p}(\delta)$, which is given by $\mathrm{sign}(\delta,\gamma)\cdot\iota_{\delta\prec\gamma}$ if $\delta$ is a face of $\gamma$, and 0 otherwise.
@@
For any tropical fan $\Sigma$ of dimension $d$, the weight function $\omega$ gives rise to a fundamental class $\Sigma,\omega]\in H_{d,d}^{BM}(\Sigma)$, and there are cap product maps $\frown[\Sigma,\omega]\colon H^{p,q}(\Sigma)\to H_{d-p,d-q}^{BM}(\Sigma)$.
@@
Let $\Sigma$ be a tropical fan with fundamental class $\Sigma,\omega]\in H_{d,d}^{BM}(\Sigma)$.
@@ 3. Tropicalizing complements of arrangements of curves
Using the above procedure, by Theorem˜2.3 the tropicalization $\operatorname{trop}(X)$ can be equipped with a fan structure consisting of one ray for each element $D$ of $\mathcal{B}$, and one two-dimensional cone between the rays for $D$ and $D^{\prime}$ if these two curves intersect in $S$. Moreover note that one may ignore the restriction on the characteristic of $K$ by Remark˜2.4. Therefore, to complete the description of the tropicalization $\operatorname{trop}(X)$, it remains to find the directions of these rays.
@@
$L_{1}$$L_{2}$$L_{3}\tkzDefPoint\tkzDefPoint\tkzDefPoint\tkzCircumCenter\tkzGetPoint\tkzDrawArc\tkzDrawArcCC$
$\widetilde{L_{1}}$$\widetilde{L_{3}}$$\widetilde{L_{2}}$$\widetilde{C}$$E_{1}$$E_{3}$$E_{2}$
Figure 1. Curve arrangement from Example˜3.1 and the compactifying divisor.
@@
Consider the arrangement of curves consisting of the coordinate axes $L_{1},L_{2},L_{3}\subset\mathbb{P}_{K}^{2}$, as well as the conic $C$ passing through the three points of intersection $L_{i}\cap L_{j}$, as shown in Figure˜1. The complement of the arrangement $X\coloneqq\mathbb{P}_{K}^{2}\setminus(L_{1}\cup L_{2}\cup L_{3}\cup C)$ is not compactified by a simple normal crossing divisor in $\mathbb{P}_{K}^{2}$, so we must blow up the three points $L_{i}\cap L_{j}$. This yields a configuration including three exceptional divisors $E_{1}$, $E_{2}$ and $E_{3}$, also displayed in Figure˜1.
@@
$0\to\langle H\rangle\oplus_{i=1}^{3}\langle e_{j}\rangle\xrightarrow{\phi}\oplus_{i=1}^{3}\langle\widetilde{L_{i}}\rangle\oplus\langle\widetilde{C}\rangle\oplus_{i=1}^{3}\langle e_{j}\rangle\xrightarrow{\pi}\mathcal{O}(X)^{*}/K^{*},$
@@ 4. Axioms for abstract arrangements of curves
In light of the description of the tropicalization of an arrangement of curves in Section˜3, we now propose an axiomatization generalizing arrangements of curves. We show that for a given arrangement, the axiomatization data extracted from the arrangement is sufficient to compute its tropicalization in its intrinsic torus.
@@ 4.1. Axioms
By the above construction, any matroid has a corresponding arroid. In particular, nonrealizable matroids are not realizable as arroids either, when requiring the specified curves to satisfy the conditions imposed by the degree function and multiplicity functions, i.e. that they must be lines intersecting in one point. However, when removing these restrictions, it may be possible to find curve arrangements having the intersection properties of the original unrealizable matroids. For nonrealizability in higher degrees, see Example˜4.6.
@@
We return to Example˜3.1. Let $\mathcal{A}=\{1,2,3,4\}$, where $i=1,2,3$ corresponds to the line $L_{i}$, and $4$ corresponds to the conic $C$. The degree function $d\colon\mathcal{A}\to\mathbb{Z}_{>0}$ is given by $d_{i}=1$ for $i<4$ and $d_{4}=2$. The multiset of points is $\mathcal{P}=\{\{1,2,4\},\{1,3,4\},\{2,3,4\}\},$ corresponding to the set of intersection points of the lines and conic. The Bézout property of $\mathcal{P}$ follows directly from the Bézout theorem for $\mathbb{P}_{K}^{2}$.
$L_{1}$$L_{2}$$L_{3}\tkzDefPoint\tkzDefPoint\tkzDefPoint\tkzCircumCenter\tkzGetPoint\tkzDrawArc\tkzDrawArcCC$
Figure 2. Curve arrangement from Example˜$4.$
@@
Consider now the generic arrangement of three lines and a conic displayed in Figure˜2. As before, let $\mathcal{A}=\{1,2,3,4\}$, where $i=1,2,3$ corresponds to the line $L_{i}$, and $4$ corresponds to the conic $C$, with the degree function $d\colon\mathcal{A}\to\mathbb{Z}_{>0}$ given by $d_{i}=1$ for $i<4$ and $d_{4}=2$. The multiset structure of the points is now more evident as
@@
In general, given an arrangement of curves $\mathcal{B}=\{C_{1},\dots,C_{n}\}$ on $\mathbb{P}_{K}^{2}$, we can create an arroid of rank two as follows. Let $\mathcal{A}\coloneqq\{1,\dots,n\}$ and the function $d$ takes $i$ to the degree of the curve $d_{i}$. For any subset $\mathbf{p}\subseteq\mathcal{B}$ such that $C_{I}=\cap_{i\in I}\;C_{i}\neq\emptyset$, let $\mathbf{p}_{I_{1}},\dots,\mathbf{p}_{I_{k}}$ be the components of $C_{I}$. For each $\mathbf{p}_{I_{j}}$, we include $\mathbf{p}$ as an element in $\mathcal{P}$ with multiplicity function $m_{\mathbf{p}}$ given by the intersection multiplicities of the curves at the component $\mathbf{p}_{I_{j}}$. Thus the point $\mathbf{p}$ may appear multiple times with potentially different multiplicity functions. Summarizing, we have show the following proposition.
@@
Figure 3. Diagrammatic representation of the nonrealizable arroid from Example˜4.6, with the “line” $L$ in dashes.
@@
Let $\mathbf{A}$ be the arroid encoding parts of the data of a (self-crossing) hexagon inscribed in a conic. Its data, most easily intuited from Figure˜3, is given by the ground set $\mathcal{A}\coloneqq\{C,1,\dots,6,L\}$, with $d(C)=2$, $d(L)=1$ and $d(i)=1$ for all $i$, the multiplicity functions are all constant taking value $1$, with multiset of points $\mathcal{P}$ consists of the points named $\mathbf{a}=\{1,3,L\}$, $\mathbf{b}=\{2,5,L\}$, $\mathbf{c}=\{4,6,L\}$, as well as the points $\{1,2,C\}$, $\{3,4,C\}$, $\{5,6,C\}$, $\{3,5,C\}$, $\{1,6,C\}$, $\{2,4,C\}$, $\{2,3\}$, $\{4,5\}$$\{1,5\}$, and $\{2,6\}$.
Pascal’s theorem, also known as the hexagrammum mysticum theorem (the reader may prefer hexagrammum nonmysticum as the name of the above arroid), states that for curves satisfying the above data, the three points $\mathbf{a}$, $\mathbf{b}$ and $\mathbf{c}$ must be colinear. By omitting $L$ from $\mathbf{c}$ in the arroid, it can no longer be realized in the projective plane. This example also generalizes the non-Pappus matroid, in light of the generalization of Pappus theorem by Pascal’s theorem, and one might construct further nonrealizable examples by working with the Cayley–Bacharach theorem.